OpenStudy (anonymous):

If you perform the experiment, "Toss Three Coins", what is the probability of getting three tails? 1/8 3/8 1/2

OpenStudy (anonymous):

well the probability of the first being tails is 1/2, the second is 1/2, and the third is 1/2 the probability of getting all three tails would then be: 1/2 x 1/2 x 1/2 = 1/8

OpenStudy (chrisasl):

It's 1/8.. 1/2 for the 1st, 2nd and 3rd>> 1/2 * 1/2 * 1/2 = 1/8